The following table shows (hypothesized) opinions voiced by two different age groups in Scotland, in response to an opinion poll on the recent referendum for Scottish independence from the U.K. Do the two Age groups differ significantly (in a statistical sense), in terms of the proportions favoring the status quo (i.e., NO to Scottish independence)? Perform a suitable statistical test to support your answer.

Age Group

Sample Size

Do NOT want Scottish independence from the U.K.

Under 30

100

48

30 - 49

150

78

use excel
Answer #1

Result:

The following table shows (hypothesized) opinions voiced by two different age groups in Scotland, in response to an opinion poll on the recent referendum for Scottish independence from the U.K. Do the two Age groups differ significantly (in a statistical sense), in terms of the proportions favoring the status quo (i.e., NO to Scottish independence)? Perform a suitable statistical test to support your answer.

Age Group

Sample Size

Do NOT want Scottish independence from the U.K.

Under 30

100

48

30 - 49

150

78

Two sample proportion test

Ho: P1 = P2   H1: P1 ≠ P2

p1=48/100=0.48

p2=78/150=0.52

P = (x1+x2)/(n1+n2)

=(48+78)/(100+150)=0.504

= -0.6197

Table value of z at 0.05 level = 1.96

Rejection Region: Reject Ho if z < -1.96 or z > 1.96

Calculated z = -0.6197 not in the rejection region

The null hypothesis is not rejected.

There is not enough evidence to conclude that proportions favoring the status quo for the two Age groups differ significantly.

Excel Addon PHStat used

Z Test for Differences in Two Proportions

Data

Hypothesized Difference

0

Level of Significance

0.05

Group 1

Number of Items of Interest

48

Sample Size

100

Group 2

Number of Items of Interest

78

Sample Size

150

Intermediate Calculations

Group 1 Proportion

0.48

Group 2 Proportion

0.52

Difference in Two Proportions

-0.04

Average Proportion

0.5040

Z Test Statistic

-0.6197

Two-Tail Test

Lower Critical Value

-1.9600

Upper Critical Value

1.9600

p-Value

0.5355

Do not reject the null hypothesis